# A Systematic Search for AGN Flares in ZTF Data Release 23

### v2

@@ 3.3 Refinement and Selection of High-Confidence Flares
In addition, we exclude sources that show signs of blending or positional confusion arising from the 3′′ cross-matching radius used in catalog construction. In particular, if multiple ZTF sources are associated with a single AGN and exhibit inconsistent separations, such as some within 1′′ and others beyond 1.5′′, we treat this as evidence of source contamination. In such cases, we extract only the light curve from sources within 1′′ and recompute the flare significance following the procedure described in Section 3.2.
@@ 4 Catalog Demographics
type Source type, including ‘Blazar’, ‘SN’, and ‘TDE’ as described in Section 4.3
ALeRCE Classification result from the ALeRCE as described in Section 4.3.
@@ 4.1 Overall Characteristics
Additionally, the flare profiles also show considerable variety, including fast-rise/slow-decay and slow-rise/fast-decay patterns (see Figure 9). The latter type appears relatively rare, and two representative light curves are presented in Figure 10. Such diversity may be associated with different physical processes, potentially linked to AMSs located in different regions of the AGN disk (liuAccretionmodifiedStarsAccretion2024).
@@ 4.2 Host AGN Types
4.3 Source Classification
We begin by selecting blazars, known for their strong variability signals. By crossmatching the flares in the AGNFCC with sources from the Roma-BZCAT Multifrequency Catalogue of Blazars  [^4] https://www.ssdc.asi.it/bzcat/ (massaro5thEditionRomaBZCAT2015) using a 3′′ matching radius, and combining them with the blazars flagged in the Million Quasars Catalog, we identify a total of 830 blazars. These sources are expected to exhibit significant variability, making them a part of our catalog.
@@ 4.3 Source Classification
We also use the Automatic Learning for the Rapid Classification of Events (ALeRCE) broker classifier (sanchez-saezAlertClassificationALeRCE2021) of the ZTF alert stream [^6] https://alerce.online/ to assign classifications to the AGN flares.
@@
5.1 Tidal Disruption Events
When a star ventures too close to a black hole (BH), the extreme tidal field can overcome the star’s self-gravity and tear it apart. The infall and accretion of the disrupted stellar material can give rise to a bright, short-lived flare known as a tidal disruption event (TDE, reesTidalDisruptionStars1988; gezariTidalDisruptionEvents2021). Although TDEs have thus far predominantly been found in inactive galaxies, TDEs are also possible to occur within the dense, gas-rich environments of AGN accretion disks with an even higher rate (karasEnhancedActivityMassive2007; chanTidalDisruptionEvents2019; mckernanStarfallHeavyRain2022; ryuInplaneTidalDisruption2024; kaurElevatedRatesTidal2025). These events may involve the disruption of main-sequence stars orbiting near the central SMBH (ryuTidalDisruptionsMainsequence2020), or occur around stellar-mass black holes embedded in AGN disks, leading to luminous signals (yangTidalDisruptionStellarmass2022).
In the AGNFCC, we find two AGN flares that match with transients classified as TDEs in the TNS: AT2020afhd and AT2022exr. Both of them were classified as TDEs based on their persistent blue optical colors, UV emission, and supporting optical spectroscopic features (hammersteinZTFTransientClassification2022; hammersteinClassificationAT2020afhdZTF20abwtifz2024). AT2020afhd exhibits a rising timescale $t_{g}\sim 20$ days and a decay timescale $t_{e}\sim 120$ days. However, it is excluded from the AGNFRC due to insufficient data coverage during the decay phase, which failed to meet our criterion 6. In contrast, AT2022exr shows similar timescales and is included in the AGNFRC. This source is particularly interesting because it exhibits repeating flare behavior, which remains challenging to fully explain with current models (langisRepeatingFlaresXray2025). The light curves of both sources are shown in Figure 13.
@@ 5.1 Tidal Disruption Events
5.2 Supernovae
Supernovae (SNe) are violent explosions caused by core collapse of massive stars (core collapse supernovae, CCSNe) or thermonuclear explosions of white dwarfs (SNe Ia). SNe can occur spatially near the galactic center or within AGN accretion disks, where the environment could potentially enhance the rate of SNe (grishinSupernovaExplosionsActive2021). For example, the rate of CCSNe in AGN disks may increase due to more active star formations in these regions (artymowiczStarTrappingMetallicity1993). On the other hand, the rate of SNe Ia could also increase with a higher frequency of double white dwarf (WD) mergers (mckernanBlackHoleNeutron2020), or by accretion onto a single WD, leading to a thermonuclear explosion where the mass is close to the Chandrasekhar limit (ostrikerViscousDragAccretion1983). Both CCSNe and SNe Ia in AGN disks have been suggested as potentially observable events, with their signatures detectable in the AGN light curves (liCorecollapseSupernovaExplosions2023; zhuThermonuclearExplosionsAccretioninduced2021).
@@ 5.2 Supernovae
5.3 Stellar mass Binary Black Hole Mergers
Stellar mass binary black hole (BBH) mergers may occur within AGN accretion disks, which provide an environment that favors both the formation and retention of massive black holes. The deep gravitational potential of AGN disks makes it difficult for merger remnants to escape (varmaEvidenceLargeRecoil2022), enabling hierarchical mergers that gradually build up BH mass (liMultimessengerHierarchicalTriple2025a). In addition, the presence of dense gas facilitates the pairing and the orbital evolution of black holes through dynamical friction and gas torques (yangHierarchicalBlackHole2019; mckernanMonteCarloSimulations2020), making AGN disks efficient for producing high-mass BBH systems (samsingAGNPotentialFactories2022).
Observationally, there is evidence that a fraction of BBH mergers detected by the LIGO-Virgo-KAGRA (LVK) collaboration may originate from AGN by investigating the spatial correlation between gravitational-wave sky maps and the position of known AGNs (zhuEvidenceFractionLIGO2025). If BBH mergers occur in AGN disks, they may produce detectable electromagnetic (EM) counterparts due to interactions between the binary (or post-merger remnant) and the surrounding medium. Possible mechanisms include ram-pressure stripping of ambient gas (mckernanRampressureStrippingKicked2019), hyper-Eddington accreting leading to Bondi explosions (wangAccretionmodifiedStarsAccretion2021), or jet launching following recoil kicks (chenElectromagneticCounterpartsPowered2024; tagawaShockCoolingBreakout2024; rodriguez-ramirezOpticalUVFlares2025; maElectromagneticFlaresAssociated2025).
grahamLightDarkSearching2023 reported seven AGN flares that may be associated with BBH mergers based on their temporal and spatial coincidence with GW events. An updated analysis by heTracingLightIdentification2025 using additional data found that only two of these flares maintain a strong correlation with GW events. Both of them are included in our AGNFRC, and their light curves are shown in Figure 15.
@@ 5.3 Stellar mass Binary Black Hole Mergers
5.4 Microlensing
Microlensing caused by the close passage of a single star or a small number of stars has been proposed as a likely mechanism for producing large amplitude variability in AGNs (lawrenceSlowblueNuclearHypervariables2016). These events are uniform in color across the rest-frame UV and optical wavelengths, and the resulting light curves display symmetric profiles. The expected characteristic timescale is $O(\mathrm{yrs})$, determined by the lens mass and the relative transverse velocity between the source and the lens (lawrenceSlowblueNuclearHypervariables2016). Under the assumption that both the background source and the lens are point-like and their relative motion is linear, the time-dependent magnification can be approximated as (grahamUnderstandingExtremeQuasar2017):
@@ 5.4 Microlensing
where $u(t)^{2}=(t-t_{0})^{2}/t_{E}^{2}+u_{0}^{2}$ is the angular distance in units of the Einstein radius between the lens and the source (for details see, e.g. schneiderGravitationalLenses1992) and $F_{\nu,obs}(t)$ and $F_{\nu,0}$ are the observed flux density and the baseline flux density.
@@
5.5 Blazars
Blazars are a subclass of AGNs characterized by relativistic jets that are closely aligned with the line of sight to the observer. This alignment results in strong Doppler boosting, producing highly variable emission across the entire electromagnetic spectrum (urryUnifiedSchemesRadioLoud1995). Their optical light curves can exhibit large-amplitude, rapid, and erratic variability, with timescales ranging from minutes to years (otero-santosOpticalVariabilityBlazar2024).
In our sample, blazars naturally account for a significant fraction of high-variability AGN flares. We find 830 blazars in the AGNFCC based on Roma-BZCAT and milliquas (see Section 4.3), representing approximately 3% of the total flares. This fraction increases to 10% in the AGNFRC, reflecting the strong variability that makes blazars more likely to pass our selection criteria. This trend is consistent with the changes shown in Figure 11 and Figure 12. Figure 17 presents two representative examples.
@@ 5.6 AGN Variability
The physical origin of such variability remains uncertain and has been debated for decades. Several theoretical mechanisms have been proposed, including the local (lyubarskiiFlickerNoiseAccretion1997) or the global accretion rate (liAccretionDiscModel2008) fluctuations, thermal fluctuations in the disk (kellyACTIVEGALACTICNUCLEUS2013), X-ray reprocessing (kubotaPhysicalModelBroadband2018), or the combination of these effects (sanchez-saezQUESTSillaAGN2018). In some cases, especially when the variability is strong or sparsely sampled, it may resemble a flare and satisfy our selection criteria.

### v3

@@ III.3 Refinement and Selection of High-Confidence Flares
In addition, we exclude sources that show signs of blending or positional confusion arising from the 3′′ cross-matching radius used in catalog construction. In particular, if multiple ZTF sources are associated with a single AGN and exhibit inconsistent separations, such as some within 1′′ and others beyond 1.5′′, we treat this as evidence of source contamination. In such cases, we extract only the light curve from sources within 1′′ and recompute the flare significance following the procedure described in Section III.2.
@@ IV Catalog Demographics
type Source type, including ‘Blazar’, ‘SN’, and ‘TDE’ as described in Section IV.3
ALeRCE Classification result from the ALeRCE as described in Section IV.3.
@@ IV.1 Overall Characteristics
Additionally, the flare profiles also show considerable variety, including fast-rise/slow-decay and slow-rise/fast-decay patterns (see Figure 9). The latter type appears relatively rare, and two representative light curves are presented in Figure 10. Such diversity may be associated with different physical processes, potentially linked to AMSs located in different regions of the AGN disk (Liu et al., 2024).
@@ IV.2 Host AGN Types
IV.3 Source Classification
We begin by selecting blazars, known for their strong variability signals. By crossmatching the flares in the AGNFCC with sources from the Roma-BZCAT Multifrequency Catalogue of Blazars  [^4] https://www.ssdc.asi.it/bzcat/ (Massaro et al., 2015) using a 3′′ matching radius, and combining them with the blazars flagged in the Million Quasars Catalog, we identify a total of 830 blazars. These sources are expected to exhibit significant variability, making them a part of our catalog.
@@ IV.3 Source Classification
We also use the Automatic Learning for the Rapid Classification of Events (ALeRCE) broker classifier (Sánchez-Sáez et al., 2021) of the ZTF alert stream [^6] https://alerce.online/ to assign classifications to the AGN flares.
@@
V.1 Tidal Disruption Events
When a star ventures too close to a black hole (BH), the extreme tidal field can overcome the star’s self-gravity and tear it apart. The infall and accretion of the disrupted stellar material can give rise to a bright, short-lived flare known as a tidal disruption event (TDE, Rees, 1988; Gezari, 2021). Although TDEs have thus far predominantly been found in inactive galaxies, TDEs are also possible to occur within the dense, gas-rich environments of AGN accretion disks with an even higher rate (Karas and Šubr, 2007; Chan et al., 2019; McKernan et al., 2022; Ryu et al., 2024; Kaur and Stone, 2025). These events may involve the disruption of main-sequence stars orbiting near the central SMBH (Ryu et al., 2020), or occur around stellar-mass black holes embedded in AGN disks, leading to luminous signals (Yang et al., 2022).
In the AGNFCC, we find two AGN flares that match with transients classified as TDEs in the TNS: AT2020afhd and AT2022exr. Both of them were classified as TDEs based on their persistent blue optical colors, UV emission, and supporting optical spectroscopic features (Hammerstein et al., 2022, 2024). AT2020afhd exhibits a rising timescale $t_{g}\sim 20$ days and a decay timescale $t_{e}\sim 120$ days. However, it is excluded from the AGNFRC due to insufficient data coverage during the decay phase, which failed to meet our criterion 6. In contrast, AT2022exr shows similar timescales and is included in the AGNFRC. This source is particularly interesting because it exhibits repeating flare behavior, which remains challenging to fully explain with current models (Langis et al., 2025). The light curves of both sources are shown in Figure 13.
@@ V.1 Tidal Disruption Events
V.2 Supernovae
Supernovae (SNe) are violent explosions caused by core collapse of massive stars (core collapse supernovae, CCSNe) or thermonuclear explosions of white dwarfs (SNe Ia). SNe can occur spatially near the galactic center or within AGN accretion disks, where the environment could potentially enhance the rate of SNe (Grishin et al., 2021). For example, the rate of CCSNe in AGN disks may increase due to more active star formations in these regions (Artymowicz et al., 1993). On the other hand, the rate of SNe Ia could also increase with a higher frequency of double white dwarf (WD) mergers (McKernan et al., 2020a), or by accretion onto a single WD, leading to a thermonuclear explosion where the mass is close to the Chandrasekhar limit (Ostriker, 1983). Both CCSNe and SNe Ia in AGN disks have been suggested as potentially observable events, with their signatures detectable in the AGN light curves (Li et al., 2023; Zhu et al., 2021).
@@ V.2 Supernovae
V.3 Stellar mass Binary Black Hole Mergers
Stellar mass binary black hole (BBH) mergers may occur within AGN accretion disks, which provide an environment that favors both the formation and retention of massive black holes. The deep gravitational potential of AGN disks makes it difficult for merger remnants to escape (Varma et al., 2022), enabling hierarchical mergers that gradually build up BH mass (Li and Fan, 2025). In addition, the presence of dense gas facilitates the pairing and the orbital evolution of black holes through dynamical friction and gas torques (Yang et al., 2019; McKernan et al., 2020b), making AGN disks efficient for producing high-mass BBH systems (Samsing et al., 2022).
Observationally, there is evidence that a fraction of BBH mergers detected by the LIGO-Virgo-KAGRA (LVK) collaboration may originate from AGN by investigating the spatial correlation between gravitational-wave sky maps and the position of known AGNs (Zhu and Chen, 2025). If BBH mergers occur in AGN disks, they may produce detectable electromagnetic (EM) counterparts due to interactions between the binary (or post-merger remnant) and the surrounding medium. Possible mechanisms include ram-pressure stripping of ambient gas (McKernan et al., 2019), hyper-Eddington accreting leading to Bondi explosions (Wang et al., 2021b), or jet launching following recoil kicks (Chen and Dai, 2024; Tagawa et al., 2024; Rodríguez-Ramírez et al., 2025; Ma et al., 2025).
Graham et al. (2023) reported seven AGN flares that may be associated with BBH mergers based on their temporal and spatial coincidence with GW events. An updated analysis by He et al. (2025) using additional data found that only two of these flares maintain a strong correlation with GW events. Both of them are included in our AGNFRC, and their light curves are shown in Figure 15.
@@ V.3 Stellar mass Binary Black Hole Mergers
V.4 Microlensing
Microlensing caused by the close passage of a single star or a small number of stars has been proposed as a likely mechanism for producing large amplitude variability in AGNs (Lawrence et al., 2016). These events are uniform in color across the rest-frame UV and optical wavelengths, and the resulting light curves display symmetric profiles. The expected characteristic timescale is $O(\mathrm{yrs})$, determined by the lens mass and the relative transverse velocity between the source and the lens (Lawrence et al., 2016). Under the assumption that both the background source and the lens are point-like and their relative motion is linear, the time-dependent magnification can be approximated as (Graham et al., 2017):
@@ V.4 Microlensing
where $u(t)^{2}=(t-t_{0})^{2}/t_{E}^{2}+u_{0}^{2}$ is the angular distance in units of the Einstein radius between the lens and the source (for details see, e.g. Schneider et al., 1992) and $F_{\nu,obs}(t)$ and $F_{\nu,0}$ are the observed flux density and the baseline flux density.
@@
V.5 Blazars
Blazars are a subclass of AGNs characterized by relativistic jets that are closely aligned with the line of sight to the observer. This alignment results in strong Doppler boosting, producing highly variable emission across the entire electromagnetic spectrum (Urry and Padovani, 1995). Their optical light curves can exhibit large-amplitude, rapid, and erratic variability, with timescales ranging from minutes to years (Otero-Santos et al., 2024).
In our sample, blazars naturally account for a significant fraction of high-variability AGN flares. We find 830 blazars in the AGNFCC based on Roma-BZCAT and milliquas (see Section IV.3), representing approximately 3% of the total flares. This fraction increases to 10% in the AGNFRC, reflecting the strong variability that makes blazars more likely to pass our selection criteria. This trend is consistent with the changes shown in Figure 11 and Figure 12. Figure 17 presents two representative examples.
@@ V.6 AGN Variability
The physical origin of such variability remains uncertain and has been debated for decades. Several theoretical mechanisms have been proposed, including the local (Lyubarskii, 1997) or the global accretion rate (Li and Cao, 2008) fluctuations, thermal fluctuations in the disk (Kelly et al., 2013), X-ray reprocessing (Kubota and Done, 2018), or the combination of these effects (Sánchez-Sáez et al., 2018). In some cases, especially when the variability is strong or sparsely sampled, it may resemble a flare and satisfy our selection criteria.